Given the following table:

Output -------- Total Cost

0 ----------------- 45

1 ----------------- 54

2 ----------------- 62

3 -----------------69

4 ----------------- 75

5 ----------------- 82

6 ----------------- 90

Part 1: The TVC of producing 4 units of output is:

Part 2: The AFC of producing 6 units of output is:  

Part 3: The marginal cost of producing the fifth unit of output is:

Part 4: The ATC of producing 5 units of output is:
Answer #1

Answer
Part 1
FC=the cost is same at all level and it is equal to total cost at Q=0
VC=TC-FC

TFC=45
TVC=TC-TFC
TVC(4)=75-45=30
--------
Part 2

AFC=FC/Q
AFC(6)=45/6=7.5
-------------
Part 3

MC(n)=(TC(n)-TC(p))/(n-p)
MC(n)=marginal cost of n th unit
TC(n)=Total cost of n units of output
TC(p)=Total cost of p unit of output
here, n>p.
MC(5)=(82-75)/(5-4)=7

-------------
Part 4
ATC=TC/Q
ATC(5)=82/5
=16.4